### v1

@@ 2.1 Satisficing Paths
This section contains several definitions useful for the of study strategic dynamics in normal-form games and presents satisficing paths as a generalization of best response paths.
Definition 2.3.
@@
The preceding definition of best response paths can be relaxed in several ways, and such relaxations are often desirable to avoid non-convergent cycling behavior. A common relaxation involves synchronizing players or incorporating inertia, so that only a subset of players switch their strategies at a given time, which can be help achieve coordination in cooperative settings (Marden and Shamma, 2012; Swenson et al., 2018a; Yongacoglu et al., 2022).
Beyond cooperative settings, the use of best response dynamics to seek Nash equilibrium may not be justified. In purely adversarial settings, for instance, best response paths cycle and fail to converge (Balcan et al., 2023), and some alternative strategic dynamics are needed to drive play to equilibrium. Consider the following generalization of the best response update:
@@
The update defined above is characterized by a “win stay, lose shift” principle, which only constrains the player to continue using a strategy when it is optimal. On the other hand, the player is not forced to use a best response when $x^{i}_{t}\notin{\rm BR}^{i}_{0}({\mathbf{x}}^{-i}_{t})$, and may experiment with suboptimal responses according to a function $f^{i}:{\mathbf{X}}\to\mathcal{X}^{i}$. [^1] As a special case, $f^{i}$ may simply be a best response selector, recovering the best response update. Allowing the function $f^{i}$ to be any function from ${\mathbf{X}}$ to $\mathcal{X}^{i}$, one generalizes best response updates and obtains a much larger set of sequences $({\mathbf{x}}_{t})_{t\geq 1}$ and greater flexibility to approach equilibrium from new directions. This motivates the following definition of satisficing paths.
Definition 2.4.
@@
Definition 2.5.
The game $\Gamma$ has the satisficing paths property if for any ${\mathbf{x}}_{1}\in{\mathbf{X}}$, there exists a satisficing path $({\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots)$ such that, for some finite $T=T({\mathbf{x}}_{1})$, the strategy profile ${\mathbf{x}}_{T}$ is a Nash equilibrium. [^2] A more general definition, involving $\epsilon\geq 0$ best responding and strategy subsets was studied in (Yongacoglu et al., 2023). In this paper, we consider true optimality and no strategic constraints, which additionally aids clarity.
Satisficing paths were initially formalized by Yongacoglu et al. (2023), who proved that two-player games and $n$-player symmetric games have the satisficing paths property. However, whether general-sum $n$-player games have the satisficing paths property was left as an open question.
3 Main result
Theorem 3.1.
@@ 3 Main result
Proof sketch. Before presenting the formal proof, we describe the intuition of its main argument. In the proof of Theorem 3.1, we construct a satisficing path from an arbitrary initial strategy ${\mathbf{x}}_{1}$ to a Nash equilibrium by repeatedly switching the strategies of unsatisfied players in a way that grows the set of unsatisfied players after the update. Once the set of unsatisfied players is maximal, we argue that a Nash equilibrium can be reached in one step by switching the strategies of the unsatisfied players. The final point represents the main technical challenge in the proof, as switching the strategies of unsatisfied players changes the objective functions for the previously satisfied players. We address this challenge by showing the existence of a Nash equilibrium on the boundary of a strategy subset in which previously satisfied players remain satisfied.
To give the complete proof, we will require some additional notation, detailed in the next subsection, and some supporting results, detailed in Appendix A and Appendix B.
3.1 Additional notation
In order to describe the construction of our satisficing path from ${\mathbf{x}}_{1}$ to a Nash equilibrium, we require the following sets, defined for any ${\mathbf{x}}\in{\mathbf{X}}$:
@@ 3.1 Additional notation
[…]
@@
3.2 Proof of Theorem 3.1
Remark 3.2.
In the proof below, we analytically construct a satisficing path from ${\mathbf{x}}_{1}$ to a Nash equilibrium. The process of selecting successive strategies ${\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots$ and switching the component strategy of each player is done centrally, by the analyst, and should not be mistaken for some type of distributed learning algorithm.
Proof 3.3.
@@
To begin, we produce a satisficing path ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{k}$ as follows. We put $t=1$, and while both $\varnothing{\rm Sat}({\mathbf{x}}_{t})\not=\varnothing$ and $\varnothing{\rm Worse}({\mathbf{x}}_{t})\not=\varnothing$, we arbitrarily fix ${\mathbf{x}}_{t+1}\in{\rm Worse}({\mathbf{x}}_{t})$ and increment $t\leftarrow t+1$. By construction, we have
[…]
@@
By the construction of the path $({\mathbf{x}}_{1},\dots,{\mathbf{x}}_{k})$, (at least) one of the following holds at index $k$: either $\varnothing{\rm Sat}({\mathbf{x}}_{k})=\varnothing$ or $\varnothing{\rm Worse}({\mathbf{x}}_{k})=\varnothing$. In other words, either no player is satisfied at ${\mathbf{x}}_{k}$, or there is no accessible strategy that grows the subset of unsatisfied players.
Case 1: $\varnothing{\rm Sat}({\mathbf{x}}_{k})=\varnothing$, and all players are unsatisfied at ${\mathbf{x}}_{k}$. In this case, we may switch the strategy of each player $i\in[n]$ to any successor strategy. That is, ${\rm Access}({\mathbf{x}}_{k})={\mathbf{X}}$. We fix an arbitrary Nash equilibrium ${\mathbf{z}}_{\star}$, put ${\mathbf{x}}_{k+1}={\mathbf{z}}_{\star}$, and let $T=k+1$. Then, $({\mathbf{x}}_{1},\dots,{\mathbf{x}}_{T})$ is a satisficing path terminating at equilibrium.
Case 2: $\varnothing{\rm Sat}({\mathbf{x}}_{k})\not=\varnothing$ and $\varnothing{\rm Worse}({\mathbf{x}}_{k})=\varnothing$. In this case, there are no accessible strategies that strictly grow the set of unsatisfied players, which is also non-empty.
Since $\varnothing{\rm Worse}({\mathbf{x}}_{k})=\varnothing$, the following holds: for any strategy ${\mathbf{y}}\in{\rm NoBetter}({\mathbf{x}}_{k})$ and any satisfied player $i\in{\rm Sat}({\mathbf{x}}_{k})$, we have that $i\in{\rm Sat}({\mathbf{y}})$. (Otherwise, if $i\in{\rm UnSat}({\mathbf{y}})$, then ${\mathbf{y}}\in{\rm Worse}({\mathbf{x}}_{k})$, since it flipped a satisfied player. But this contradicts the emptiness of ${\rm Worse}({\mathbf{x}}_{k})$.)
@@
From the set containment in (2), it is clear that ${\mathbf{x}}_{\star}\notin{\rm NoBetter}({\mathbf{x}}_{k})$, since ${\rm NoBetter}({\mathbf{x}}_{k})$ consists of strategies accessible from ${\mathbf{x}}_{k}$ in which unsatisfied agents remain unsatisfied, while the previously unsatisfied agents are satisfied at ${\mathbf{x}}_{\star}$. We now state a key technical lemma, which asserts that although ${\mathbf{x}}_{\star}$ does not belong to ${\rm NoBetter}({\mathbf{x}}_{k})$, it does lie on its boundary.
Lemma 3.4.
There exists a sequence $\{{\mathbf{y}}\}_{t=1}^{\infty}$, with ${\mathbf{y}}_{t}\in{\rm NoBetter}({\mathbf{x}}_{k})$ for each $t$, such that
$\lim_{t\to\infty}{\mathbf{y}}_{t}={\mathbf{x}}_{\star}.$
A proof of Lemma 3.4 given in Appendix B.
To conclude the proof, we will argue that ${\mathbf{x}}_{\star}$ is in fact a Nash equilibrium for the game $\Gamma$. That is, in addition to (2), we also have ${\rm Sat}({\mathbf{x}}_{k})\subset{\rm Sat}({\mathbf{x}}_{\star})$. To do so, we introduce functions $F^{i}:{\mathbf{X}}\to\mathbb{R}$ for each player $i\in[n]$, given by
$F^{i}(x^{i},{\mathbf{x}}^{-i})=\max_{a^{i}\in\mathbb{A}^{i}}R^{i}(\delta_{a^{i}},{\mathbf{x}}^{-i})-R^{i}(x^{i},{\mathbf{x}}^{-i}),$
for each ${\mathbf{x}}=(x^{i},{\mathbf{x}}^{-i})\in{\mathbf{X}}$. The functions $\{F^{i}\}_{i=1}^{n}$ have the following useful properties, which are well known (Maschler et al., 2020), and are summarized in Appendix A. For each player $i\in[n]$,
  1. a.
    $F^{i}$ is continuous on ${\mathbf{X}}$,
  2. b.
    $F^{i}({\mathbf{x}})\geq 0$ for all ${\mathbf{x}}\in{\mathbf{X}}$, and
  3. c.
Let $({\mathbf{y}}_{t})_{t=1}^{\infty}$ be a sequence in ${\rm NoBetter}({\mathbf{x}}_{k})$ converging to ${\mathbf{x}}_{\star}$, which exists by Lemma 3.4. For any previously satisfied player $i\in{\rm Sat}({\mathbf{x}}_{k})$, since $\varnothing{\rm Worse}({\mathbf{x}}_{k})=\varnothing$ and ${\mathbf{y}}_{t}\in{\rm NoBetter}({\mathbf{x}}_{k})$, from a previous observation, we have that $i\in{\rm Sat}({\mathbf{y}}_{t})$. Equivalently, $x^{i}_{k}\in{\rm BR}^{i}_{0}({\mathbf{y}}^{-i}_{t})$. Re-writing this using the function $F^{i}$ and the notation $y^{i}_{t}=x^{i}_{k}$ for satisfied players $i\in{\rm Sat}({\mathbf{x}}_{k})$, we have
$F^{i}(y^{i}_{t},{\mathbf{y}}^{-i}_{t})=0,\quad\forall t\in\mathbb{N},$
for any $i\in{\rm Sat}({\mathbf{x}}_{k})$. By continuity of $F^{i}$, we have
@@
establishing that player $i$ is satisfied at ${\mathbf{x}}_{\star}$. Since $i\in{\rm Sat}({\mathbf{x}}_{k})$ was generic, we have ${\rm Sat}({\mathbf{x}}_{k})\subset{\rm Sat}({\mathbf{x}}_{\star})$. Then, by (2), we also had ${\rm UnSat}({\mathbf{x}}_{k})\subset{\rm Sat}({\mathbf{x}}_{\star})$, so indeed ${\rm Sat}({\mathbf{x}}_{\star})=[n]$, and ${\mathbf{x}}_{\star}$ is a Nash equilibrium accessible from ${\mathbf{x}}_{k}$.
We put $T=k+1$ and ${\mathbf{x}}_{T}={\mathbf{x}}_{\star}$, which completes the proof, since $({\mathbf{x}}_{1},\dots,{\mathbf{x}}_{T})$ is a satisficing path terminating at a Nash equilibrium.
@@ 4 Discussion
Multi-agent reinforcement learning algorithms based on the “win stay, lose shift” principle characteristic of satisficing paths are especially well suited to decentralized applications, since players are often able to estimate the performance of their current strategy as well as the performance of an optimal strategy, even under partial information. In decentralized problems, coordinated search of the set ${\mathbf{X}}$ of strategy profiles for a Nash equilibrium is typically infeasible, and players must select successor strategies in a way the depends only on quantities that can be locally accessed or estimated.
For instance, consider a trivial coordinated search method, where player $i$ selects $x^{i}_{t+1}$ uniformly at random from $\mathcal{X}^{i}$ whenever ${\mathbf{x}}_{t}$ was not a Nash equilibrium and selects $x^{i}_{t+1}=x^{i}_{t}$ only when ${\mathbf{x}}_{t}$ is a Nash equilibrium. This process is clearly ill-suited to decentralized applications, because player $i$’s strategy update depends on both a locally estimable condition (whether player $i$ is best responding to ${\mathbf{x}}^{-i}_{t}$) as well as a condition that cannot be locally estimated (whether another player $j\not=i$ is best responding to ${\mathbf{x}}^{-j}_{t}$.) The satisfaction (“win stay”) constraint plays a key role as a local stopping condition for satisficing paths, and rules out coordinated search of the set ${\mathbf{X}}$ such as the trivial update outlined above. Examples of decentralized or partially decentralized learning algorithms leveraging satisficing paths in their analysis include (Foster and Young, 2006; Marden et al., 2009; Arslan and Yüksel, 2017; Yongacoglu et al., 2023). The analytic results of this paper suggest that algorithms such as these can be extended to wider classes of games and enjoy equilibrium guarantees under different informational constraints on the players.

### v2

@@ 2.1 Satisficing Paths
We now present the concept of satisficing paths as generalized best response paths.
Definition 3.
@@
The preceding definition of best response paths can be relaxed in several ways, and such relaxations are often desirable to avoid non-convergent cycling behavior (see [37] for an example). A common relaxation involves synchronizing players or incorporating inertia, so that only a subset of players switch their strategies at a given time, which can be help achieve coordination in cooperative settings [32, 48, 51]. Beyond cooperative settings, the use of best response dynamics to seek Nash equilibrium may not be justified. In purely adversarial settings, for instance, best response paths cycle and fail to converge [3], and some alternative strategic dynamics are needed to drive play to equilibrium. Consider the following generalization of the best response update:
@@
The update defined above is characterized by a “win–stay, lose–shift" principle [11, 44], which only constrains the player to continue using a strategy when it is optimal. On the other hand, the player is not forced to use a best response when $x^{i}_{t}\notin{\rm BR}^{i}_{0}({\mathbf{x}}^{-i}_{t})$, and may experiment with suboptimal responses according to a function $f^{i}:{\mathbf{X}}\to\mathcal{X}^{i}$. [^1] As a special case, $f^{i}$ may simply be a best response selector, recovering the best response update. Allowing the function $f^{i}$ to be any function from ${\mathbf{X}}$ to $\mathcal{X}^{i}$, one generalizes best response updates and obtains a much larger set of sequences $({\mathbf{x}}_{t})_{t\geq 1}$ and greater flexibility to approach equilibrium from new directions. This motivates the following definition of satisficing paths.
Definition 4.
@@
Definition 5.
The game $\Gamma$ has the satisficing paths property if for any ${\mathbf{x}}_{1}\in{\mathbf{X}}$, there exists a satisficing path $({\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots)$ such that, for some finite $T=T({\mathbf{x}}_{1})$, the strategy profile ${\mathbf{x}}_{T}$ is a Nash equilibrium. [^2] A more general definition, involving $\epsilon\geq 0$ best responding and strategy subsets was studied in [52]. In this paper, we consider true optimality and no strategic constraints, which additionally aids clarity.
Satisficing paths were initially formalized in [52], where it was proved that two-player games and $n$-player symmetric games have the satisficing paths property. However, whether general-sum $n$-player games have the satisficing paths property was left as an open question. We answer this open question in Theorem 1, presented in the next section.
3 Existence of paths in normal-form games
Theorem 1.
@@ 3 Existence of paths in normal-form games
Proof sketch. Before presenting the formal proof, we describe the intuition of its main argument. In the proof of Theorem 1, we construct a satisficing path from an arbitrary initial strategy ${\mathbf{x}}_{1}$ to a Nash equilibrium by repeatedly switching the strategies of unsatisfied players in a way that grows the set of unsatisfied players after the update. Once the set of unsatisfied players is maximal, we argue that a Nash equilibrium can be reached in one step by switching the strategies of the unsatisfied players. The final point represents the main technical challenge in the proof, as switching the strategies of unsatisfied players changes the objective functions for the previously satisfied players. We address this challenge by showing the existence of a Nash equilibrium on the boundary of a strategy subset in which previously satisfied players remain satisfied.
To give the complete proof, we will require some additional notation, detailed below, and some supporting results, detailed in Appendix A and Appendix B.
Additional notation. We require notation for the following sets, defined for any ${\mathbf{x}}\in{\mathbf{X}}$:
@@
${\rm Worse}({\mathbf{x}}):=$$\left\{{\mathbf{y}}\in{\rm NoBetter}({\mathbf{x}}):{\rm UnSat}({\mathbf{x}})\subsetneq{\rm UnSat}({\mathbf{y}})\right\}$
@@
3.1 Proof of Theorem 1
Remark 1.
In the proof below, we analytically construct a path from ${\mathbf{x}}_{1}$ to a Nash equilibrium. The process of selecting strategies ${\mathbf{x}}_{1},{\mathbf{x}}_{2},\cdots$ and switching the component strategy of each player is done centrally, by the analyst, and should not be interpreted as a learning algorithm.
Proof.
@@ 3.1 Proof of Theorem 1
To begin, we produce a satisficing path ${\mathbf{x}}_{1},\dots,{\mathbf{x}}_{k}$ as follows. We put $t=1$, and while both ${\rm Sat}({\mathbf{x}}_{t})\not=\varnothing$ and ${\rm Worse}({\mathbf{x}}_{t})\not=\varnothing$, we arbitrarily fix ${\mathbf{x}}_{t+1}\in{\rm Worse}({\mathbf{x}}_{t})$ and increment $t\leftarrow t+1$. By construction, we have
${\rm UnSat}({\mathbf{x}}_{1})\subsetneq\cdots\subsetneq{\rm UnSat}({\mathbf{x}}_{t})\subsetneq{\rm UnSat}({\mathbf{x}}_{t+1})$
@@
By the construction of the path $({\mathbf{x}}_{1},\dots,{\mathbf{x}}_{k})$, (at least) one of the following holds at index $k$: either ${\rm Sat}({\mathbf{x}}_{k})=\varnothing$ or ${\rm Worse}({\mathbf{x}}_{k})=\varnothing$. In other words, either no player is satisfied at ${\mathbf{x}}_{k}$, or there is no accessible strategy that grows the subset of unsatisfied players.
Case 1: ${\rm Sat}({\mathbf{x}}_{k})=\varnothing$, and all players are unsatisfied at ${\mathbf{x}}_{k}$. In this case, we may switch the strategy of each player $i\in[n]$ to any successor strategy. That is, ${\rm Access}({\mathbf{x}}_{k})={\mathbf{X}}$. We fix an arbitrary Nash equilibrium ${\mathbf{z}}_{\star}$, put ${\mathbf{x}}_{k+1}={\mathbf{z}}_{\star}$, and let $T=k+1$. Then, $({\mathbf{x}}_{1},\dots,{\mathbf{x}}_{T})$ is a satisficing path terminating at equilibrium.
Case 2: ${\rm Sat}({\mathbf{x}}_{k})\not=\varnothing$ and ${\rm Worse}({\mathbf{x}}_{k})=\varnothing$. In this case, there are no accessible strategies that strictly grow the set of unsatisfied players.
Since ${\rm Worse}({\mathbf{x}}_{k})=\varnothing$, the following holds: for any strategy ${\mathbf{y}}\in{\rm NoBetter}({\mathbf{x}}_{k})$ and any satisfied player $i\in{\rm Sat}({\mathbf{x}}_{k})$, we have that $i\in{\rm Sat}({\mathbf{y}})$. (Otherwise, if $i\in{\rm UnSat}({\mathbf{y}})$, then ${\mathbf{y}}\in{\rm Worse}({\mathbf{x}}_{k})$, since it flipped a satisfied player. But this contradicts the emptiness of ${\rm Worse}({\mathbf{x}}_{k})$.)
@@
From (2), it is clear that ${\mathbf{x}}_{\star}\notin{\rm NoBetter}({\mathbf{x}}_{k})$, since ${\rm NoBetter}({\mathbf{x}}_{k})$ consists of strategies accessible from ${\mathbf{x}}_{k}$ in which unsatisfied agents remain unsatisfied, while the previously unsatisfied agents are satisfied at ${\mathbf{x}}_{\star}$. We now state a key technical lemma, which asserts that although ${\mathbf{x}}_{\star}$ does not belong to ${\rm NoBetter}({\mathbf{x}}_{k})$, it is a limit point of this set. A proof of Lemma 1 given in Appendix B.
Lemma 1.
If ${\rm Worse}({\mathbf{x}}_{k})=\varnothing$, then there exists a sequence $\{{\mathbf{y}}\}_{t=1}^{\infty}$, with ${\mathbf{y}}_{t}\in{\rm NoBetter}({\mathbf{x}}_{k})$ for each $t$, such that $\lim_{t\to\infty}{\mathbf{y}}_{t}={\mathbf{x}}_{\star}.$
We will argue that ${\mathbf{x}}_{\star}$ is a Nash equilibrium for the original game $\Gamma$. For each player $i\in[n]$, we introduce a function $F^{i}:{\mathbf{X}}\to\mathbb{R}$ given by $F^{i}(x^{i},{\mathbf{x}}^{-i})=\max_{a^{i}\in\mathbb{A}^{i}}R^{i}(\delta_{a^{i}},{\mathbf{x}}^{-i})-R^{i}(x^{i},{\mathbf{x}}^{-i}),$ for each ${\mathbf{x}}=(x^{i},{\mathbf{x}}^{-i})\in{\mathbf{X}}$. The functions $\{F^{i}\}_{i=1}^{n}$ have the following useful properties, which are well known [35], and are summarized in Appendix A. For each player $i\in[n]$: (a) $F^{i}$ is continuous on ${\mathbf{X}}$; (b) $F^{i}({\mathbf{x}})\geq 0$ for all ${\mathbf{x}}\in{\mathbf{X}}$; (c) for any ${\mathbf{x}}^{-i}\in{\mathbf{X}}^{-i}$, a strategy $x^{i}$ is a best response to ${\mathbf{x}}^{-i}$ if and only if $F^{i}(x^{i},{\mathbf{x}}^{-i})=0$.
Let $({\mathbf{y}}_{t})_{t=1}^{\infty}$ be a sequence in ${\rm NoBetter}({\mathbf{x}}_{k})$ converging to ${\mathbf{x}}_{\star}$, which exists by Lemma 1. For any previously satisfied player $i\in{\rm Sat}({\mathbf{x}}_{k})$, since ${\rm Worse}({\mathbf{x}}_{k})=\varnothing$ and ${\mathbf{y}}_{t}\in{\rm NoBetter}({\mathbf{x}}_{k})$, from a previous observation, we have that $i\in{\rm Sat}({\mathbf{y}}_{t})$. Equivalently, $x^{i}_{k}\in{\rm BR}^{i}_{0}({\mathbf{y}}^{-i}_{t})$. Re-writing this using the function $F^{i}$ and the notation $y^{i}_{t}=x^{i}_{k}$ for satisfied players $i\in{\rm Sat}({\mathbf{x}}_{k})$, we have $F^{i}(y^{i}_{t},{\mathbf{y}}^{-i}_{t})=0$ for all $t\in\mathbb{N}$ and for any $i\in{\rm Sat}({\mathbf{x}}_{k})$. By continuity of $F^{i}$, we have
@@
establishing that player $i$ is satisfied at ${\mathbf{x}}_{\star}$, and thus that ${\rm Sat}({\mathbf{x}}_{k})\subset{\rm Sat}({\mathbf{x}}_{\star})$. Then, by (2), we had ${\rm UnSat}({\mathbf{x}}_{k})\subset{\rm Sat}({\mathbf{x}}_{\star})$, hence ${\rm Sat}({\mathbf{x}}_{\star})=[n]$, and ${\mathbf{x}}_{\star}$ is a Nash equilibrium accessible from ${\mathbf{x}}_{k}$. We put $T=k+1$ and ${\mathbf{x}}_{T}={\mathbf{x}}_{\star}$, which completes the proof, since $({\mathbf{x}}_{1},\dots,{\mathbf{x}}_{T})$ is a satisficing path terminating at a Nash equilibrium. ∎
3.2 Algorithmic insights from the proof of Theorem 1
When coupled with a MARL algorithm that uses an exploratory satisficing strategy update, play will be driven along satisficing paths. Theorem 1 shows that for any starting strategy profile, some such path connects the strategy profile to an equilibrium, and so a sufficiently exploratory strategy update may drive play to equilibrium along a satisficing path. This offers important insights for the design of MARL algorithms. The first takeaway from Theorem 1 is that play can be driven to equilibrium by changing only the strategies of those players who are not best responding. In particular, this means that a satisfied agent does not need to continue updating its strategy after it becomes satisfied. As we will discuss in the next section, this property is helpful in distributed and decentralized multi-agent systems, where agents are able to assess whether they are satisfied but may not be able to assess whether the overall system is at equilibrium.
A second, more subtle takeaway comes from the proof of Theorem 1 and relates to the unorthodox and counterintuitive exploration scheme used to drive play to equilibrium. In the proof, one sees that suboptimal—and perhaps even reward-deteriorating—strategic updates were key to driving play to equilibrium along a satisficing path. As we outline below, this construction runs against the conventional approaches to designing MARL algorithms, and it can be used to avoid common pitfalls of MARL algorithms such as cyclical behavior.
At a high level, many existing multi-agent learning algorithms update the strategy parameter in a reward-improving direction at each step. A related approach, described earlier, increments the strategy parameter in a regret-minimizing direction, which has a similar effect. While such algorithms are sensible from the point of view of a single self-interested individual, they may fail to drive play to a Nash equilibrium when all players adopt similar algorithms [36, 18, 37]. To address this non-convergence issue, one recurring algorithmic modification involves manipulating step sizes, either with a mixture of fast agents and slow agents [13] or with each individual varying its step sizes according to its performance [8]. However, such approaches only come with provable convergence guarantees in select subclasses of games with exploitable structure. In instances where step size manipulation does not (or cannot) yield convergence, the analysis of Theorem 1 may offer an alternative route to algorithm modification.
With these two takeaways in mind, we envision at least two design principles that will be useful for future MARL algorithms. First, strategic updating may incorporate some measure of randomness when a player is not satisfied. This principle has been previously used with some success, but comes with a drawback relating to complexity. A second principle, which we believe to be new, leverages the second takeaway above, involving counterintuitive path construction: players may alternate between reward-improving periods (during which strategy updates are done in a conventional way that improves the agent’s reward) and suboptimal periods (during which reward-deteriorating and/or random strategy updates may be used). The timing of such periods or the extent of the randomness in strategic updates may be made to depend on whether cycles in the strategy iterates were detected. By incorporating suboptimal exploration in an adaptive manner, a MARL algorithm can break cycles as needed but rely on conventional algorithms the remainder of the time.
@@ 4 Discussion
Extension to Markov games
This paper focused on normal-form games with finitely many actions per player due to the central position that normal-form games occupy in game theory. Indeed, insights and intuition developed in normal-form games are helpful for understanding more complex models of strategic interaction. Of special note, finite normal-form games can be generalized to model dynamic strategic environments where rewards and environmental parameters evolve over time according to the history of play. We now describe the extension of Theorem 1 to Markov games, one generalization of finite normal-form games that is a popular model in MARL. Due to space limitations, a formal model for Markov games is postponed to Appendix C.
In an $n$-player Markov game, agents interact across discrete time. Each agent $i\in[n]$ observes a sequence of state variables $\{s_{t}\}_{t\geq 1}$ taking values in a finite state space $SS\SS$ and selects a sequence of actions $\{a^{i}_{t}\}_{t\geq 1}$ taking values in a finite action set $\mathbb{A}^{i}$. In this dynamic model, player $i$’s reward in period $t$, denoted $r^{i}_{t}=r^{i}(s_{t},{\mathbf{a}}_{t})$, depends on both the action profile ${\mathbf{a}}_{t}$ and also on the state $s_{t}$. The state process evolves according to a (jointly controlled) transition probability function $\mathcal{T}$ as $s_{t+1}\sim\mathcal{T}(\cdot|s_{t},{\mathbf{a}}_{t})$. Rewards are discounted across time using a discount factor $\gamma\in(0,1)$, and player $i$ attempts to maximize its expected $\gamma$-discounted return. In this generalization of finite normal-form games, policies (defined as mappings from states to probability distributions over actions) generalize mixed strategies, and the solution concept of Markov perfect equilibrium refines the concept of Nash equilibrium and serves as a popular stability objective for MARL algorithm designers [53].
Partial results for multi-state Markov games have previously been obtained in special classes of games and used to produce MARL algorithms [52]. The analysis presented in this paper uses a rather different approach that seems promising for extending those results. In the proof of Theorem 1, we used functions $\{F^{i}\}_{i=1}^{n}$ to characterize best responding in a finite normal-form game. In fact, analogous functions can also be obtained for policies in multi-state Markov games, and these functions satisfy the same desired properties invoked in the proof of Theorem 1 (c.f. [52, Lemmas 2.10-2.13]). For this reason, and due to the central role of continuity in our proof, it seems likely that Theorem 1 can be extended to general-sum Markov games. However, one aspect of the extension remains open, namely the generalization of Lemma 1. In Appendix C, we describe the issue that precludes direct generalization of our normal-form proof of Lemma 1, but we note that this appears to be related only to the proof technique rather than a fundamental obstacle to the generalization.
@@
Multi-agent reinforcement learning algorithms based on the “win–stay, lose–shift” principle characteristic of satisficing paths are especially well suited to decentralized applications, since players are often able to estimate the performance of their current strategy as well as the performance of an optimal strategy, even under partial information. In decentralized problems, coordinated search of the set ${\mathbf{X}}$ of strategy profiles for a Nash equilibrium is typically infeasible, and players must select successor strategies in a way the depends only on quantities that can be locally accessed or estimated.
For instance, consider a trivial coordinated search method, where player $i$ selects $x^{i}_{t+1}$ uniformly at random from $\mathcal{X}^{i}$ whenever ${\mathbf{x}}_{t}$ was not a Nash equilibrium and selects $x^{i}_{t+1}=x^{i}_{t}$ only when ${\mathbf{x}}_{t}$ is a Nash equilibrium. This process is clearly ill suited to decentralized applications, because player $i$’s strategy update depends on both a locally estimable condition (whether player $i$ is best responding to ${\mathbf{x}}^{-i}_{t}$) as well as a condition that cannot be locally estimated (whether another player $j\not=i$ is best responding to ${\mathbf{x}}^{-j}_{t}$.) The satisfaction (win–stay) constraint plays a key role as a local stopping condition for satisficing paths, and rules out coordinated search of the set ${\mathbf{X}}$ such as the trivial update outlined above. Examples of decentralized or partially decentralized learning algorithms leveraging satisficing paths in their analysis include [19, 33, 1, 52]. The analytic results of this paper suggest that algorithms such as these can be extended to wider classes of games and enjoy equilibrium guarantees under different informational constraints on the players.